OpenStudy (ttp):

Solve the equation Seven plus n plus one equals five times (one minus n) for n. –0.5 0.5 1.5 2.0

OpenStudy (anonymous):

Seven plus n plus one equals five times (one minus n) 7 + n + 1 = 5 (1 - n) 8 + n = 5 -5n |-n 8 = 5 - 6n |+6n -8 6n = 5-8 6n = -3 |:6 n = -0.5